OpenStudy (anonymous):

what is the number one half of one quater of one tenth of 400

OpenStudy (mokeira):

\[\frac{ 1 }{2 }\times \frac{ 1 }{ 4 }\times \frac{ 1 }{ 10 }\times 400=simplify\]

OpenStudy (lxelle):

First divide 400 by 10 400/10 = 40 then divide the result by 4 40/4 = 10 and lastly divide the result by 2 10/2 = ??

OpenStudy (anonymous):

tnx dear